### v1

# Simulating nationwide coupled disease and fear spread in an agent-based model

Abstract
Human behavior and disease spread evolve together over the course of any disease outbreak. We present a dynamic model of that relationship based on the coupled spread of a disease and fear of that disease, and implement this model in a simulator, EpiCast, that simulates disease spread between individuals in a realistic synthetic population. In our model, fear of the disease spreads through both in-person contact and broadcast media, driving people to adopt behaviors that reduce disease spread. In order to better understand the dynamics of our model, we create and compare a range of compartmental surrogate models to analyze the impact of including various disease states. We also compare how Epicast-simulated outbreaks unfold across a number of scenarios with different types of behavioral responses to fear and a range of different parameter values. We find that the addition of asymptomatic, exposed, and pre-symptomatic disease states can impact both the rate at which an outbreak progresses and its overall trajectory. In addition, the combination of non-local fear spread through broadcasters and strong preventative measures by afraid individuals generally produces multiple epidemic waves, which only occur for a narrow range of parameter values with purely local fear spread.
keywords:
Agent-based model,Coupled contagions,Infectious diseases,Disease propagation1 Introduction
Looking back on the early stages of the COVID-19 pandemic reveals a dizzying array of responses across all levels of society. From state-mandated lockdowns and school closures to individual masking, testing, and vaccination, these responses shaped not just personal experiences of the pandemic, but observed outcomes at a societal level as well [1, 2, 3, 4].
Unfortunately, accurately anticipating individual behavior has proved challenging for the similarly diverse array of models developed to simulate the disease’s spread. The most common method of accounting for behavioral changes has been to use exogenous data sources, such as cell phone mobility data or historical public health policies, to impose top-down behaviors [5]. However, exogenous approaches struggle to capture the two-way feedback between behavioral changes and disease spread, since they mostly treat behavior as an input parameter, able to be scheduled in advance [6]. Truly dynamic behaviors require incorporating a behavioral model into the heart of a simulation, able to respond flexibly to changes in the global or local simulation state.
However, most such endogenous behavioral models operate within high-level compartmental models where a series of ordinary differential equations (ODEs) govern the changing proportion of the population in different disease states [7]. Such models can only answer population-level questions about disease spread because they smooth over the heterogeneity of real-world populations. In contrast, agent-based models (ABMs) capture this heterogeneity by directly modeling the disease state and behavior of each individual in the simulated population. This capability has enabled the use of ABMs for retrospective analysis of COVID-19 spread patterns [8], evaluation of public health interventions [9, 10, 11], and the study of differential COVID-19 outcomes across regions and demographic groups [12]. However, ABMs rarely include endogenous behaviors [7], and the few existing ABMs which do only simulate relatively small populations consisting of at most a million people [13].
To address these gaps, we present a coupled contagion model of interdependent disease and fear spread implemented in EpiCast. EpiCast is a production ABM capable of simulating the spread of an infectious disease – such as COVID-19 – on a surrogate of the full US population while accounting for varying demographic and occupational distributions, as well as commute and flight patterns across the country [9, 14]. In our model, fear of a disease spreads alongside the disease itself, influencing individuals to adopt various protective behaviors that either reduce their number of contacts or the likelihood they are infected by any given contact. This fear can spread through either in-person contact – much like the disease – or through broadcast media. To our knowledge, this is the first work to model the coupled dynamics of two contagions in an agent-based simulation of this scale and complexity.
Additionally, we present a series of ODE-based prototype models we developed to bridge the gap between existing coupled contagion models and our implementation in EpiCast. EpiCast includes several disease states beyond the standard susceptible, (symptomatic) infectious, recovered (SIR) model, such as asymptomatic and presymptomatic infectious states. As most prior coupled contagion models have been based on the SIR model, implementing our model in EpiCast required adapting existing ODE-based models to account for these new disease states before translating the resulting system into the agent-based context. We then analyzed how the incremental inclusion of disease states in these models changes their dynamics.
Our key contributions include:
1.
The formulation of an ODE model of coupled disease and fear spread which accounts for asymptomatic disease transmission.
2.
The design and implementation of an agent-based analogue of this coupled contagion model in a production simulation, incorporating both local and non-local fear spread.
3.
A sensitivity analysis of the impact of several key model parameters which govern the behavioral response to fear on salient features of the simulation results, including the total attack rate and the emergence of multiple epidemic waves.
2 Related work
Endogenous behavior models incorporate a framework for human behavior directly into a simulation in a way that allows the simulated population to react flexibly to novel scenarios. Endogenous models have proven able to produce multiple epidemic waves and at times dramatically improve model predictions [15, 16], and generally fall into three categories [7]:
1.
Feedback loop models, where behavior directly depends on disease prevalence. One noteworthy example by Rahmandad et al. outperformed the majority of the U.S. Center for Disease Control COVID-19 Forecasting Hub models by scaling the transmission rate in a SEIR compartmental model by a prevalence-dependent scaling factor [15]. Kasa et al. provide an overview of such models [17].
2.
Game theoretic models, where individuals select an option that maximizes a utility function describing their preferred outcomes. Huang et al. survey models of this type [18].
3.
Coupled contagion models, where an individual’s behaviors depends on their beliefs, which are in turn informed by a combination of their neighbor’s beliefs and information regarding the overall state of the pandemic, either of which can spread in a manner similar to the disease itself. Epstein et al. present such a model where two contagions – fear of the disease and the disease itself – spread simultaneously through a population [19], adding a third contagion – fear of vaccination – in a later version of the model [20].
Several models also combine elements from multiple categories above. Poletti et al. developed a behavioral model for H1N1 Influenza where individuals choose between protective and normal behaviors so as to maximize a utility function based on the recent disease prevalence [21]. Jovanovic et al. compare models of vaccine uptake using an objective function which incorporates population-level prevalence information and local-level information spread through neighbors in a information diffusion network [22].
2.1 Agent-based modeling of endogenous behavior
Despite the range of developed behavioral models, the vast majority are built on top of standard compartmental models, most of which use a simple feedback loop. In particular, there are only a handful of agent based models (ABMs) which use coupled contagion models [7]. Epstein et al. develop a relatively simple ABM where 1800 agents move around a 2D grid and can spread fear and disease to agents in adjacent cells, and can either flee, isolate or retain the same behavior while fearful. Note that while they demonstrate the emergence of waves in the ordinary differential equation (ODE) version of their model, they fail to do so for the ABM [19]. Rajabi et al. also use a grid-based ABM, but only have agents quarantine when afraid. They demonstrate how multiple waves can emerge in their simulation and simulate the impact of travel restrictions and contact tracing on disease spread in a 5000 agent population [23]. Palomo-Briones et al. present a more complex but still spatially explicit 400 agent model where agents have a short visit schedule consisting of visits to their home, social activities, and economic activities, and can chose to wear masks, social distance, or wash hands based on a combination of other agents’ beliefs and the results of their previous behavior [24]. Curiel et al. use a small 5000-agent network-based ABM where a disease spreads alongside anti-vaccine beliefs under different network topologies and vaccination strategies [25]. Mao presents a detailed simulation of the Buffalo metropolitan area with nearly a million agents, with a schedule of activities drawn from demographic data. Agents can decide to use flu prophylaxis using a threshold-based decision model based on either infection risk or prophylaxis adoption among contacts [13]. While these models range in complexity, only Rajabi et al., with the simplest model, demonstrate an ability to produce multiple epidemic waves. In addition, the largest simulated population is smaller than most U.S. states and tuned for a narrow geographic area, which limits the usefulness of its results for state-level and national policy-makers.
2.2 The EpiCast simulator
In stark contrast to the smaller scale of the coupled contagion ABMS described above lie HPC-scale simulators such as EpiCast. EpiCast simulates disease spread on a massive artificial surrogate population of the full U.S. containing over 322 million agents, simulated people with individualized behavior [9, 14]. In the baseline version of EpiCast, this behavior relies on a combination of fixed exogenous responses based on demographics and top-down public health interventions taken as input. The population used in EpiCast is built using UrbanPop [26], a framework which uses American Community Survey (ACS) Data to create high-fidelity synthetic populations. Compared to similar frameworks, UrbanPop captures more of the heterogeneity in real-world populations, producing cross-sectional representations of agents across sociodemographic, economic, educational, and mobility characteristics. UrbanPop assigns each member of the generated population to a specific household in a specific census block group, in a manner which ensures the aggregate demographics of the generated block group correspond roughly to those observed in census data. In EpiCast, these households form the basis for dividing the population into communities, within which agents have varying contact intensities defined by whether they share a household, neighborhood, workplace, school, etc. These contact intensities determine the pairwise likelihood that disease transmission occurs between two agents within the same community. Agents then commute between a home and a work or school community each day in a manner consistent with Department of Transportation commute flow data, and occasionally engage in long distance travel to geographically distant communities based on flight data.
3 Methods
When designing our model, we set out to adapt an existing coupled contagion ODE model to function within EpiCast, which requires accounting for all of the disease states used in EpiCast. We first construct a SIR-based ODE-model as a simple baseline, then incrementally incorporate new disease states until we have accounted for all those present in the original EpiCast simulation. Once this design is complete, we then modify the model to function in the individualized context of EpiCast. In the process, we introduce both short- and medium-distance mechanisms for fear spread and two types of behavioral responses to fear with an eye toward reproducing the twin epidemic waves found in the ODE models.
3.1 Coupled contagion ODE models
We take a streamlined version of Epistein et al.’s model of coupled fear and disease spread citeepstein2008coupled as our starting point due to its flexibility and ability to produce multiple epidemic waves. We then incrementally add the asymptomatic, presymptomatic and exposed disease states present in EpiCast to this model. For brevity, we refer to this series of models using the Cartesian product of the two types of states. For example, we would refer to a standard model paired with a fear model with Neutral and Afraid states as an coupled model.
Over the course of this process, we designed the following series of models:
1.
, the closest to Epstein et al.’s original model. The flow rate from neutral and afraid compartments is slower by a factor of from the Recovered state. corresponds to the Epstein model.
2.
, distinguishing between symptomatic and asymptomatic Infectious and Recovered states. Of the disease states, only has a reduced flow rate from neutral to afraid compartments.
3.
, adds an Exposed state to represent the incubation period and a Presymptomatic state to allow for disease spread prior to the potential onset of symptoms. This is the closest to the model implemented in EpiCast.
For brevity, we present only the system in full. The flow diagram is shown in Fig S1 and the full system of equations is given in Section S1.
3.2 Coupled contagion agent-based model
The EpiCast coupled contagion model can be thought of as a translation of the model into the individualized context of EpiCast. This requires reworking aspects of the basic functionality of the original EpiCast simulator to facilitate fear spread and tie an individual’s behavior to their fear state.
3.2.1 Modeling fear spread
We implement two methods of fear spread in EpiCast: one which operates based on an agent’s local neighborhood of contacts in a manner similar to disease spread, and a second that approximates the effect of traditional broadcast media.
In order to implement local fear spread, we took advantage of EpiCast’s preexisting methodology for computing disease spread, refactoring the code to allow for fear spread using different per-contact transmission probabilities. Fear is spread by symptomatic and afraid agents and countered by agents who know they have recovered from an infection and have a neutral fear state (i.e. agents in disease state and fear state ).
For non-local fear spread, we created broadcasters capable of taking one of three positions on the disease. They may either: (1) spread fear, (2) counter fear spread, or (3) remain neutral. In order to site these broadcasters, we begin by examining the occupations of agents in each community, which are identified by north American Industry Classification System (NAICS) code in EpiCast. If we identify at least one agent with a relevant NAICS code in a given community we locate a broadcaster there and assign all agents with the relevant occupations to work there. Each broadcaster then is able to potentially influence anyone within the census tract in which they are located, and starts by taking a neutral position.
Once initialized, a broadcaster’s position is determined by a combination of the fear states of its employees and the level of disease spread in the broadcast area. When the proportion of employees at a broadcaster who are afraid of the disease crosses a certain threshold, , the broadcaster begins spreading fear. We then record the current infection rate, the number of new infections in the broadcast area as a proportion of its population, as . In future timesteps, the broadcaster determines its position by comparing the current infection rate, , against as follows:
1.
If , the broadcaster begins countering fear spread.
2.
Otherwise, if , the broadcaster takes a neutral position.
3.
Otherwise, the broadcaster spread fear of the disease.
where .
For each agent in the broadcast area of a given broadcaster, they have a fixed chance, , of being influenced by that broadcaster in a given timestep if the broadcast is not neutral and their fear state does not correspond to the broadcaster’s position.
3.2.2 Modeling behavior
We modeled two different ways in which agents may change their behavior when they are afraid: (1) taking protective measures (such as masking and hand-washing) while maintaining their normal routine, and (2) withdrawing from their normal routine and remaining at home. We model this first measure by reducing the susceptibility of agents in state , similar to how this is handled in [20] and in the ODE model.
In the second case, we modify an agent’s routine so that they remain in their home community, neither commuting to their work community nor engaging in long distance travel, and only interacting with other members of their household. If currently traveling, they immediately return to their home community. Note that this is different from a complete quarantine, as they can still transmit the disease or their fear to their household. If agents are afraid, they can either withdraw purely due to fear with a fixed probability for a given duration , or due to having a symptomatic infection (i.e. being in state ). Note that any agent with a symptomatic infection will withdraw if their symptoms are severe enough for them to be hospitalized, regardless of their fear state.
4 Results
We designed our experiments in order to answer three main research questions:
RQ 1
How does the introduction of additional disease compartments change outbreak dynamics?
RQ 2
How does the addition of different types of withdrawal and fear spread impact the results of simulations conducted in EpiCast?
RQ 3
How do key parameters in our models relate to important simulation characteristics, such as the total attack rate or the number of epidemic waves?
We describe below how we designed experiments to answer these questions, then outline their results.
4.1 Experimental setup
First, we sought to compare solutions to each ODE model with an initial value problem (IVP) solver in order to identify scenarios under which each of our ODE models produce multiple epidemic waves, providing a comparison of when the same parameter values produce substantially different outcomes between models. This was intended to directly address RQ 1 and provide insights into which parameters might be of interest for RQ 3. Parameter values are shown in Table S1.
Second, we sought to compare several scenarios with different types of withdrawals – for hospitalized agents (hospitalized withdrawals), those who are symptomatic and afraid (sick withdrawals), and those who are merely afraid (fear-only withdrawals) – and both purely local (based on physical contact with agents in the same community) and non-local (based on broadcasters operating within a given census tract) fear spread modes in order to address RQ 2. For these scenarios, we use a synthetic population of about 322 million agents representing the 48 contiguous U.S. states and Washington DC. We seeded index cases on a per-county basis using estimates of active cases on 25 March 2020 derived from data from The New York Times, based on reports from state and local health agencies. For the first five days of the simulation, for an average of 856 thousand seed cases per day (about 0.265% of the population). Parameter values are shown in Table S2.
Finally, we investigated RQ 3 by conducting a series of sensitivity analyses comparing EpiCast outputs for a range of values of and , which determine the strength of an agent’s behavioral response to fear of the disease. We drew these values from the range , trying five values for each parameter. We ran each cell twice – once with only local fear spread and once with both local and broadcaster-based fear spread. These experiments were conducted on the contiguous U.S. dataset described above. Choices for parameters other than and correspond to Scenarios (c) and (d) from Table S2 (Or, equivalently, Scenarios (e) and (f)) without and with broadcasters, respectively. In order to compute the number of epidemic waves programmatically, we first compute a centered 7-day rolling maximum of new case counts – so as to avoid identifying peaks caused solely by weekly variation – then call the find_peaks function from the Scipy signal package (v1.14.0) in Python with a prominence of . This value indicates that new cases should fall by at least 1 new infection per 10,000 people before rising again if we are to consider a given local maximum to be the peak of a given epidemic wave. We found this value to avoid spurious peaks (i.e. those in the long tail of some outbreaks) while still capturing those most visually prominent in the data.
4.2 Comparing ODE models
For the ODE comparison experiments, Fig. 1 shows the solutions corresponding the parameter choices that either failed to produce a second epidemic wave within 360 days (top) or produced such a wave (bottom). For the model, we observe that changing the likelihood that those who recovered from a symptomatic infection become newly afraid of the disease, , has a dramatic impact on whether or not a second wave occurs. With the proportion of the population which are susceptible or afraid largely plateau around 0.8 and 0.6, respectively, after 360 days, as shown in Fig. 1(a). In contrast, these proportions plateau at around 0.35 and 0, respectively, when we use , as shown in Fig. 1(d). In additional, a second wave only occurs in the latter case, peaking a bit after 150 days.
When we introduce asymptomatic disease states, as in the model, the situation changes further. Keeping all parameters used in the run the same with this model results in 80% of the population still susceptible and 40% afraid after 360 days, and no second epidemic wave, as shown in Fig. 1(b). Changing the susceptibility of afraid agents, , from 25% to 35%, however, is sufficient to produce a second wave in this model, albeit a smaller one which peaks much later at around 275 days, as shown in Fig. 1(e). Under this scenario, about half of the population is still susceptible after 360 days although this proportion is still decreasing. The proportion afraid, while approaching zero by 360 days, does so about 160 days later than in the two wave scenario without asymptomatic infections (Fig. 1(d)).
Introducing an exposed state and a pre-symptomatic infectious state into the model produces the model, and further changes in the solution space. Using the same pair of parameter values as with the model results in faster disease spread and decay in fear levels. Fig. 1(c) shows that about 70% and 20% of the population is susceptible and afraid, respectively, after 360 days with in this model. Additionally, while we do not see a second wave peak within 360 days here, we do see the start of such a wave and the lack of a plateau of fear levels that corresponds to a single-wave outbreak in the other modes. When using with the model, Fig. 1(f) shows that fear levels reach zero and a second epidemic peak as with the model. However, both these events happen much sooner in the former model, although still about 100 days later than in the model.
Altogether, with respect to RQ 1 we find that adding asymptomatic states to the ODE models significantly delays the progression of multiple waves and reduces the rate of infections and fear decay as well as the total attack rate, while adding an exposed state has the opposite effect. Neither addition significantly changes the peak of fear levels. As for RQ 3, we find that changing the values of and both showed the potential to change both whether or not we observe a second wave and the total attack rate.
4.3 Comparing selected scenarios in EpiCast
We now examine the results of the various scenarios evaluated for the coupled contagion model implemented in EpiCast, so as to explore RQ 2. Fig. 2(a) showcases the differences in the count, height, and timing of epidemic peaks between scenarios. In particular, we see the highest peak in the scenario with only hospitalization-based withdrawals (hosp), as well as the highest total attack rate. This is followed by the scenario where we allow withdrawals based on having symptoms (hosp+sick) and pure-fear withdrawals from fear as well (hosp+sick+fear), which also show a single peak. The scenario with the smallest peak (and also the lowest total attack rate) comes from using hospitalization-based and symptomatic withdrawals along with a susceptibility reduction from fear (hosp+sick+reduced_sus), which also has a single peak. The only scenarios wherein we observe multiple peaks are those where we add broadcasters as a fear spread mechanism with purely fear-based withdrawals (hosp+sick+fear+bc)and a fear-based susceptibility reduction (hosp+sick+reduced_sus+bc), with the former having higher peaks than the latter, and the second peak occurring at roughly the same time in both cases.
We observe that fear levels peak after 35 to 60 days at around 60-70% of the population in most cases, with the exception of the pure-fear withdrawal scenario without broadcasters (hosp+sick+fear), where fear levels peak after 85 days at around 38% of the population. In most cases, fear levels fall dramatically from this peak before the end of the simulation. The scenarios with reduced susceptibility due to fear both buck this trend, with fear levels only falling slightly by day 200 without broadcasters (hosp+sick+reduced_sus) and dropping slowly by around 15 percentage points in total with broadcasters (hosp+sick+reduced_sus+bc). We note that, in the absence of broadcasters, the reduced susceptibility case (orange dashed line in Fig. 2(b)) displays the highest fear levels for most of the simulation and the pure-fear withdrawal case the lowest (red dashed line in Fig. 2(b)). In the presence of broadcasters, fear trends more closely resemble other runs. Fear levels fall significantly after their peak in the susceptibility reduction case (blue line in Fig. 2(b)), when they remaining roughly constant in the absence of broadcasters. Conversely, fear levels rose much higher before falling in the pure-fear withdrawal case (green dotted line in Fig. 2(b)), peaking at above 60% of the population rather than less than 40%.
In both scenarios with broadcasters, more broadcasters spread fear than countered it throughout the majority of the run, though in the pure-fear withdrawal case this changed shortly before the end of the run. Far more broadcasters were spreading fear in the susceptibility reduction scenario than in the pure-fear withdrawal scenario for the entire duration (compare blue and green lines in Fig. 2(c)), while the reverse is true for fear countering for most of the duration of the run (Fig. 2(d)). Broadcasters were not used in the other scenarios.
Fig. 3 highlights the geographical distribution of new cases for selected time steps and US states for two of the scenarios shown above. Fig. 3(a) and 3(b) represent the results of the scenario with pure-fear withdrawals from fear with purely local fear spread, where Fig. 3(c) and 3(d) do so when fear spread through broadcasters is additionally allowed (hosp+sick+fear and hosp+sick+fear+bc, respectively, in Fig. 2). The time steps chosen represent peaks (days 50, 100, and 130) and a trough (day 75) in the latter, multi-wave scenario. Particular US states were chosen to represent different epidemic trajectories.
For the former case, Fig. 3(b) shows how the state-level trajectories differ primarily based on the height of their peak, though Texas (TX) does generally lag behind the trends displayed by the other selected states by a handful of days, with the exception of Colorado (CO) in the declining phase of the outbreak.
In contrast, we see significant divergence between states in the broadcaster scenario, seen in Fig. 3(d). South Dakota (SD) and Washington (WA) both have high initial peaks which transition into lower secondary peaks, with WA having an earlier first peak and consistently lower case counts throughout. Utah (UT), Colorado (CO), and Texas (TX), by contrast, each have low initial peaks and higher secondary peaks. For UT, there is very little decline in cases before the upward trend towards the second peak begins, with case counts peaking roughly 20-30 days before the other states. UT and CO reach similar heights to each other for each peak, while TX has a lower second peak, only slightly above WA.
4.4 EpiCast sensitivity analysis
In order to get a better sense of how varying the behavioral response to fear within EpiCast influences the overall trajectory of disease outbreaks, as per RQ 3, we ran a number of simulation varying two influential parameters. The first was the likelihood that an agent who is afraid of the disease – but not currently showing symptoms – withdraws from their regular schedule of activities purely due to fear, . The second was the scaling factor applied to an afraid agent’s susceptibility to infection, , which represents the impact of afraid agents complying with NPIs that do not directly change their daily schedule, such as masking, maintaining social distance while in public, and hand-washing.
The results of conducting this parameter sweep with only local fear spread are shown in Fig. 4. Fig. 4(a) displays the overall trajectories of each outbreak, with the proportion of new cases shown in orange, that of afraid agents show in blue, and epidemic peaks indicated by the dashed vertical lines. Note that all parameter combinations produce a single epidemic wave. The simulation which comes the closest corresponds to and . Further experiments, shown in Fig. 4(b), revealed that a very narrow range of values – , and – produce multiple waves, but that a deviation by as little in either parameter can be sufficient to return to the single wave scheme.
Fig. 4(c) shows the total attack rate of each simulated outbreak.
We observe the highest attack rate with and , which represents a minimal behavioral response to fear (and corresponds to Scenario (b) from Section 4.3), followed by the cases with with lower values of . Notably, fear levels are much lower in the latter case. The lowest attack rates, in contrast, are observed when and are both small.
Higher values of correspond to fewer afraid agents and result reduced variation in fear trajectories for different values of . For , we observe that fear levels rapidly plateau for and rise to a peak before falling to varying degrees for . This drop in fear levels is more pronounced for greater values of , when the overall attack rate is greater.
The results of conducting this parameter sweep with both local and broadcaster-based fear spread are shown in Fig. 5. The overall trajectories of each outbreak are shown in Fig. 5(a) with the dashed vertical lines indicating where peaks occur, and the orange and blue lines indicating proportion of new cases and afraid agents, respectively. Fig. 5(b) highlights which combinations of parameters produce one and two epidemic waves. All single wave scenarios occur with low values of , and either a high or low value of (i.e the upper right corner of Fig. 5(b) and the top two cells in the left of that figure). All other parameter combinations produce two waves. These two waves are generally about the same height with but the second is both taller and longer than the first in most other cases.
The same parameter combinations that produce a single wave also produce the most extreme attack rates, as highlighted by Fig. 5(c). The upper right corner (with high values of and low values of ) contains cells with the highest total attack rates, while the single wave runs in the upper left corner exhibit the lowest. Compared to the corresponding simulations without broadcaster-based fear spread, those with broadcaster-spread fear tend to show more moderate attack rates. For example, attack rates fall from (bottom row of Fig. 4(c)) to (bottom row of Fig. 5(c)) when broadcasters are introduced with . Conversely, attack rates largely rise in the top half of the first two columns the figure, with for instance the attack rate with increasing from 1.4% to 3.9%, and that with rising from .
Fear levels, in turn, generally follow the same basic trajectory, rising to an initial peak before falling in first an initial steep decline, then a secondary slow decline. The transition between these two periods of decline generally corresponds to when cases begin rising for a second epidemic wave. The main exceptions are found in the top left and right corners of Fig. 5(a), a plateau after a very short decline from the initial peak in the former case and a single accelerating decline for two cells in the latter. These mostly correspond to the cells with a single epidemic wave. The height of the initial peaks decreases as increases and fear levels fall farther when more cases occur.
5 Discussion
Examining the fear levels in each scenario, shown in Fig. 2(b), offers some insights into why we might observe multiple peaks primarily in cases with non-local fear spread via broadcasters. In cases where fear reduces agent susceptibility to infection, in the absence of fear spread through broadcasters (hosp+sick+reduced_sus) fear levels never fall far below their peak around 50 days. This is likely due to infection levels never rising to the point where there is a sufficient number of agents in the state to cause afraid agents to lose their fear. This is in contrast to what occurs with fear spread through broadcasters (hosp+sick+reduced_sus+bc), where fear levels quickly fall after reaching their peak. This suggests that the influence of broadcasters is sufficient to lower fear levels to the point where the disease can start spreading again, causing a second wave. Roughly the opposite appears to occur in the cases where we allow purely fear-based withdrawals; without the influence of broadcasters (hosp+sick+fear), fear levels remain lower than in any other scenario for the duration of the run. This is likely because withdrawals reduce both fear and disease spread in in this mode there is a high chance that any given afraid person will withdraw. Conversely, with the influence of broadcasters (hosp+sick+fear+bc) fear levels rise much higher before falling, though they remain lower than in all other scenarios except that with only hospitalization-based withdrawals. This suggests broadcasters allow fear to spread even when most fearful agents remain withdrawn.
These explanations are supported by examining the number of broadcasters spreading (see Fig. 2(c)) and countering fear spread (see Fig. 2(d)) over time. The period from about 25-50 where fear levels peak and then begin to fall in the reduced susceptibility scenario with broadcasters is also the period when we see the number of broadcasters countering fear spread dramatically increase before reaching a much lower sustained growth rate which again corresponds to a slow linear decrease in the number of afraid agents. The number of fear spreading broadcasters follows a similar trend to fear levels, here, albeit with a smaller falloff after the peak. Similarly, we see overall fear levels and the number of fear-spreading broadcasters in the pure-fear withdrawal case, with the dramatic reduction in fearful agents between days 50 and 100 matched by decline in fear-spreading broadcasters of similar magnitude.
In addition, the very narrow range of parameter values for which we do observe multiple waves with purely local fear spread appears to fall in the midst of a phase transition between a regime of small, short outbreaks and large, long outbreaks as represented by the cell , and the cell in Fig. 4(a). The former regime corresponds to lower values of which represent a strong NPI response to fear, implying that fear-induced behavior reduces disease spread sufficiently to forestall a larger outbreak here. The latter regime, by contrast, represents minimal behavioral responses to fear, as seen by the fact it centers on , where fear has no impact of the behavior of agents without symptoms. Intriguingly, these two regimes also appear with broadcaster-based fear spread, and correspond to the only cells that do not exhibit multiple waves in Fig. 5(a).
Altogether, these observations suggest that we observe two waves for opposite reasons in these cases. With reduced susceptibility, fear levels generally do not fall far enough for disease transmission to pick back up and produce a second wave without broadcasters countering fear spread. Conversely, with pure-fear fear-based withdrawals, fear levels generally fail to rise high enough fast enough for the first wave to leave a large reservoir of susceptible agents without broadcasters increasing the rate of fear spread. In either case, non-local fear spread, as represented by broadcasters, appears to facilitate a transition from high initial fear levels to low final fear levels in the presence of fear-based behaviors that significantly reduce disease spread. We posit that this transition only occurs under a few narrow range of conditions in the case of purely local fear spread as a result of fear both spreading and receding relatively slowly. In contrast, the faster rate at which fear levels rise and fall in the presence of broadcasters may explain why multiple epidemic waves are more commonly observed under these conditions. In other words, in order to produce multiple epidemic waves in this model, fear must first spread significantly faster than the disease and then recede faster than it as well. Within this framing, it is clear why broadcasters facilitate the formation of multiple waves: information (and the behavioral changes it engenders) both spreads and is forgotten more rapidly when it is less tightly constrained by physical distance.
The results we present in this paper are perhaps best understood as an exploration of the high-level implications of our model. Though the foundational elements of Epicast are mature and have been calibrated extensively, there remain significant obstacles to doing so for the fear-based components of the model. While some survey data is available describing various behaviors throughout part of the pandemic [27], it is more difficult to tie specific behaviors to specific influences or rationales. We thus confine ourselves to presenting the model itself and exploring the impact of varying parameter values, and leave calibration efforts to future work.
In addition, we acknowledge that the parameter sweep based sensitivity analysis study presented above is necessarily incomplete. While we explore the full possible range of parameter values for and , we do so at a relatively coarse resolution, so as to limit the total number of simulation runs. As evidenced by the multi-wave scenarios presented in Fig. 4(b), this coarseness means that this study may miss dynamics that only emerge for a small range of parameter values. In addition, there are many simulation parameters beyond the two we chose to study. Extending the parameter sweep method to the other model parameters would be prohibitively expensive, but in the absence of such an experiment it is unclear how varying those parameters might affect the dynamics in Epicast.
6 Conclusion
In this work we presented a series of models which capture the coupled spread of fear of a disease and the disease itself. We do so with several ODE models in order to demonstrate the impact of adding additional disease states to the model, representing asymptomatic, exposed, and pre-symptomatic disease states, and an ABM covering all disease states with both local and non-local mechanisms for fear spread implemented.
We find that the total attack rate is reduced by the addition of asymptomatic states and increased by the addition of a pre-symptomatic state in our ODE models, and the reverse is true for the rate at which an outbreak progresses.
In our ABM implementation in EpiCast, we primarily observe multiple epidemic waves when using our non-local fear spread mechanism, which represents traditional broadcast media. We find that a combination of (1) high initial fear levels, (2) a strong behavioral response to fear, and (3) low fear levels later in the simulation produces multiple waves in our model. Thus, one potential explanation for why we only observe multiple waves with purely local fear spread for a very small range of parameter choices is that fear generally spreads (and recedes) much more slowly in the absence of broadcasters in our model.
Acknowledgments
This work was supported by the U.S. Department of Energy, Office of Science, Office of Advanced Scientific Computing Research and by cooperative agreement CDC-RFA-FT-23-0069 from the CDC’s Center for Forecasting and Outbreak Analytics. Its contents are solely the responsibility of the authors and do not necessarily represent the official views of the Centers for Disease Control and Prevention.
This work was performed at Los Alamos National Laboratory (LANL), an equal opportunity employer, which is operated by Triad National Security, LLC, for the National Nuclear Security Administration (NNSA) of the U.S. Department of Energy (DOE) under contract #19FED1916814CKC. The funders had no role in study design, data collection and analysis, decision to publish, or preparation of the manuscript. The findings and conclusions in this report are those of the authors and do not necessarily represent the official position of LANL. This research used resources provided by the Darwin testbed at LANL which is funded by the Computational Systems and Software Environments subprogram of LANL’s Advanced Simulation and Computing program (NNSA/DOE).
This material is based upon work supported by the U.S. Department of Energy, Office of Science, Office of Advanced Scientific Computing Research, Department of Energy Computational Science Graduate Fellowship under Award Number DE-SC0021
Author contributions statement
J.K., P.C.A, S.Y.V., and T.C.G. developed the conceptual framework and J.K. wrote the majority of the code for the behavioral models presented in this work. This code included the ODE models, and the extension to EpiCast, with P.C.A. contributing some code to the later. P.C.A. and T.C.G. provided insights into the original design and implementation of EpiCast which informed the implementation of the behavioral model. J.K., A.B., S.Y.V., and T.C.G. conceived the experiments, and J.K. conducted them and analyzed the results. J.K. prepared the initial draft of the manuscript, along with all figures. All authors reviewed the manuscript.
Data availability statement
We intend to release all output data from the simulations described in this paper available via Zenodo link upon publication, along with all scripts necessary to generate the plots contained within this manuscript from those results. These scripts are unfortunately still in the process of being reviewed for release by the project sponsors and are thus not available at this time.
Competing interests
The authors declare no competing interests.
Supplementary Material
S1 Coupled contagions ODE models
We provide the full formulation of the system of ODEs which define the model below. This system is summarized in the flow diagram in Fig. S1. Let be the size of the compartment with disease state and fear state . The system is as follows:
S2 Model parameters
Parameter values used in ODE model comparison experiments (Section 4.2) are shown in Table S1 while those used in the EpiCast scenario comparision experiments (Section 4.3) are shown in Table S2.
Variable | Description | Experiment | ||||||||
(a) | (b) | (c) | (d) | (e) | (f) | |||||
disease transmission rate | 0.2 | 0.2 | 0.2 | 0.2 | 0.2 | 0.2 | ||||
disease recovery rate | 0.5 | 0.5 | 0.5 | 0.5 | 0.5 | 0.5 | ||||
0.6 | 0.6 | 0.6 | 0.6 | ||||||
0.4 | 0.4 | 0.4 | 0.4 | ||||||
0.75 | 0.75 | 0.75 | 0.75 | ||||||
0.5 | 0.5 | ||||||||
0.22 | 0.22 | 0.22 | 0.22 | 0.22 | 0.22 | ||||
baseline fear loss rate | 0.05 | 0.05 | 0.05 | 0.05 | 0.05 | 0.05 | ||||
fear loss contact rate | 0.44 | 0.44 | 0.44 | 0.44 | 0.44 | 0.44 | ||||
1 | 0 | 0 | 0 | 0 | 0 | ||||
1 | 1 | 1 | 1 | 1 | 1 | ||||
0.25 | 0.25 | 0.25 | 0.35 | 0.25 | 0.35 |
Variable | Description | Scenario | ||||||||
(a) | (b) | (c) | (d) | (e) | (f) | |||||
disease transmission rate | 0.2 | 0.2 | 0.2 | 0.2 | 0.2 | 0.2 | ||||
disease recovery rate | 0.5 | 0.5 | 0.5 | 0.5 | 0.5 | 0.5 | ||||
0.6 | 0.6 | 0.6 | 0.6 | 0.6 | 0.6 | ||||
0.4 | 0.4 | 0.4 | 0.4 | 0.4 | 0.4 | ||||
0.75 | 0.75 | 0.75 | 0.75 | 0.75 | 0.75 | ||||
0.5 | 0.5 | 0.5 | 0.5 | 0.5 | 0.5 | ||||
0.22 | 0.22 | 0.22 | 0.22 | 0.22 | 0.22 | ||||
baseline fear loss rate | 0.05 | 0.05 | 0.05 | 0.05 | 0.05 | 0.05 | ||||
fear loss contact rate | 0.44 | 0.44 | 0.44 | 0.44 | 0.44 | 0.44 | ||||
0 | 0 | 0 | 0 | 0 | 0 | ||||
1 | 1 | 1 | 1 | 1 | 1 | ||||
1 | 1 | 1 | 1 | 0.35 | 0.35 | ||||
0 | 1 | 1 | 1 | 1 | 1 | ||||
0 | 0 | 0.65 | 0.65 | 0 | 0 | ||||
0 | 0 | 0 | 0.25 | 0 | 0.25 | ||||
0.5 | 0.5 | ||||||||
0.25 | 0.25 | ||||||||
0.125 | 0.125 |

### v2

@@
-Simulating nationwide coupled disease and fear spread in an agent-based model
+Simulating nationwide coupled disease and fear spread in an agent-based model Joy Kitson
 Abstract
-Human behavior and disease spread evolve together over the course of any disease outbreak.
-We present a dynamic model of that relationship based on the coupled spread of a disease and fear of that disease, and implement this model in a simulator, EpiCast, that simulates disease spread between individuals in a realistic synthetic population.
-In our model, fear of the disease spreads through both in-person contact and broadcast media, driving people to adopt behaviors that reduce disease spread.
-In order to better understand the dynamics of our model, we create and compare a range of compartmental surrogate models to analyze the impact of including various disease states.
-We also compare how Epicast-simulated outbreaks unfold across a number of scenarios with different types of behavioral responses to fear and a range of different parameter values.
-We find that the addition of asymptomatic, exposed, and pre-symptomatic disease states can impact both the rate at which an outbreak progresses and its overall trajectory.
-In addition, the combination of non-local fear spread through broadcasters and strong preventative measures by afraid individuals generally produces multiple epidemic waves, which only occur for a narrow range of parameter values with purely local fear spread.
-keywords:
-Agent-based model,Coupled contagions,Infectious diseases,Disease propagation1 Introduction
-Looking back on the early stages of the COVID-19 pandemic reveals a dizzying array of responses across all levels of society.
-From state-mandated lockdowns and school closures to individual masking, testing, and vaccination, these responses shaped not just personal experiences of the pandemic, but observed outcomes at a societal level as well [1, 2, 3, 4].
-Unfortunately, accurately anticipating individual behavior has proved challenging for the similarly diverse array of models developed to simulate the disease’s spread.
+Human cognitive responses, behavioral responses, and disease dynamics co-evolve over the course of any disease outbreak, and can result in complex feedbacks.
+We present a dynamic agent-based model that explicitly couples the spread of disease with the spread of fear surrounding the disease, implemented within the EpiCast simulation framework.
+EpiCast models transmission across a realistic synthetic population, capturing individual-level interactions.
+In our model, fear propagates through both in-person contact and broadcast media, prompting individuals to adopt protective behaviors that reduce disease spread.
+In order to better understand these coupled dynamics, we create and compare a range of compartmental surrogate models to analyze the impact of including various disease states.
+Additionally, we compare a range of behavioral scenarios within EpiCast, varying the level and intensity of fear and behavioral change.
+Our results show that the addition of asymptomatic, exposed, and pre-symptomatic disease states can impact both the rate at which an outbreak progresses and its overall trajectory.
+Moreover, the combination of non-local fear spread via broadcasters and strong behavioral responses by fearful individuals generally leads to multiple epidemic waves, an outcome that occurs only within a narrow parameter range when fear spreads purely through local contact.
+Accounting for the coupled spread of fear and disease is critical for understanding disease dynamics and designing timely, targeted responses to emerging infectious threats.
+Looking back on the early stages of the COVID-19 pandemic reveals a dizzying array of responses across all levels of society, across the entire world.
+The response to the COVID-19 pandemic may have been the largest coordinated behavior change in the history of humanity.
+From state-mandated lockdowns and school closures to individual masking, testing, and vaccination, these responses shaped both personal experiences of the pandemic, and also observed outcomes at a societal level [1, 2, 3, 4].
+Unfortunately, accurately anticipating individual behavior has proved challenging for the similarly diverse array of models developed to simulate the disease’s spread [5].
 The most common method of accounting for behavioral changes has been to use exogenous data sources, such as cell phone mobility data or historical public health policies, to impose top-down behaviors [6].
@@
 Truly dynamic behaviors require incorporating a behavioral model into the heart of a simulation, able to respond flexibly to changes in the global or local simulation state.
-However, most such endogenous behavioral models operate within high-level compartmental models where a series of ordinary differential equations (ODEs) govern the changing proportion of the population in different disease states [7].
+However, most such endogenous behavioral models operate within high-level compartmental models where a series of ordinary differential equations (ODEs) govern the changing proportion of the population in different disease states [8, 9].
 Such models can only answer population-level questions about disease spread because they smooth over the heterogeneity of real-world populations.
@@
 3.
-Coupled contagion models, where an individual’s behaviors depends on their beliefs, which are in turn informed by a combination of their neighbor’s beliefs and information regarding the overall state of the pandemic, either of which can spread in a manner similar to the disease itself.
+Coupled contagion models, where an individual’s behaviors depends on their beliefs.
+An individual’s beliefs are in turn informed by a combination of their neighbor’s beliefs and information regarding the overall state of the pandemic, either of which can spread in a manner similar to the disease.
 Epstein et al.
@@
 Epstein et al.
-develop a relatively simple ABM where 1800 agents move around a 2D grid and can spread fear and disease to agents in adjacent cells, and can either flee, isolate or retain the same behavior while fearful.
-Note that while they demonstrate the emergence of waves in the ordinary differential equation (ODE) version of their model, they fail to do so for the ABM [19].
+[21] develop a relatively simple ABM where 1800 agents, simulated people with individualized behavior, move around a 2D grid.
+Agents can spread fear and disease to agents in adjacent cells, and can either flee, isolate or retain the same behavior while fearful.
+Note that while they demonstrate the emergence of waves in the ordinary differential equation (ODE) version of their model, they fail to do so for the ABM.
 Rajabi et al.
-also use a grid-based ABM, but only have agents quarantine when afraid.
-They demonstrate how multiple waves can emerge in their simulation and simulate the impact of travel restrictions and contact tracing on disease spread in a 5000 agent population [23].
+[25] also use a grid-based ABM, but only have agents quarantine when fearful.
+They demonstrate how multiple waves can emerge in their simulation and simulate the impact of travel restrictions and contact tracing on disease spread in a 5000 agent population.
 Palomo-Briones et al.
-present a more complex but still spatially explicit 400 agent model where agents have a short visit schedule consisting of visits to their home, social activities, and economic activities, and can chose to wear masks, social distance, or wash hands based on a combination of other agents’ beliefs and the results of their previous behavior [24].
+[26] present a more complex but still spatially explicit 400 agent model where agents have a short visit schedule consisting of visits to their home, social activities, and economic activities, and can chose to wear masks, social distance, or wash hands based on a combination of other agents’ beliefs and the results of their previous behavior.
 Curiel et al.
-use a small 5000-agent network-based ABM where a disease spreads alongside anti-vaccine beliefs under different network topologies and vaccination strategies [25].
-Mao presents a detailed simulation of the Buffalo metropolitan area with nearly a million agents, with a schedule of activities drawn from demographic data.
-Agents can decide to use flu prophylaxis using a threshold-based decision model based on either infection risk or prophylaxis adoption among contacts [13].
-While these models range in complexity, only Rajabi et al., with the simplest model, demonstrate an ability to produce multiple epidemic waves.
+[27] use a small 5000-agent network-based ABM where a disease spreads alongside anti-vaccine beliefs under different network topologies and vaccination strategies.
+Mao [15] presents a detailed simulation of the Buffalo metropolitan area with nearly a million agents, with a schedule of activities drawn from demographic data.
+Agents can decide to use flu prophylaxis using a threshold-based decision model based on either infection risk or prophylaxis adoption among contacts.
+While these models range in complexity, only Rajabi et al., with the simplest model, demonstrates an ability to produce multiple epidemic waves.
 In addition, the largest simulated population is smaller than most U.S.
@@
 2.2 The EpiCast simulator
-In stark contrast to the smaller scale of the coupled contagion ABMS described above lie HPC-scale simulators such as EpiCast.
+The smaller scales of the coupled contagion ABMs described above contrast sharply with that of HPC-scale simulators like EpiCast.
 EpiCast simulates disease spread on a massive artificial surrogate population of the full U.S.
-containing over 322 million agents, simulated people with individualized behavior [9, 14].
+containing over 322 million agents [11, 16].
 In the baseline version of EpiCast, this behavior relies on a combination of fixed exogenous responses based on demographics and top-down public health interventions taken as input.
@@
 Compared to similar frameworks, UrbanPop captures more of the heterogeneity in real-world populations, producing cross-sectional representations of agents across sociodemographic, economic, educational, and mobility characteristics.
-UrbanPop assigns each member of the generated population to a specific household in a specific census block group, in a manner which ensures the aggregate demographics of the generated block group correspond roughly to those observed in census data.
+UrbanPop assigns each member of the generated population to a specific household in a specific census block group, in a manner which ensures the aggregate demographics of the generated block group correspond precisely to those observed in census data.
 In EpiCast, these households form the basis for dividing the population into communities, within which agents have varying contact intensities defined by whether they share a household, neighborhood, workplace, school, etc.
@@
 3 Methods
-When designing our model, we set out to adapt an existing coupled contagion ODE model to function within EpiCast, which requires accounting for all of the disease states used in EpiCast.
-We first construct a SIR-based ODE-model as a simple baseline, then incrementally incorporate new disease states until we have accounted for all those present in the original EpiCast simulation.
-Once this design is complete, we then modify the model to function in the individualized context of EpiCast.
+In designing our model, we seek to adapt an existing coupled contagion ODE model to function within EpiCast, which requires accounting for all of the disease states present in the original code.
+We first construct a SIR-based ODE-model as a simple baseline, then incrementally incorporate new disease states until we have accounted for all those used by EpiCast.
+Once this design is complete, we then modify the model to function in an individualized context.
 In the process, we introduce both short- and medium-distance mechanisms for fear spread and two types of behavioral responses to fear with an eye toward reproducing the twin epidemic waves found in the ODE models.
 3.1 Coupled contagion ODE models
-We take a streamlined version of Epistein et al.’s model of coupled fear and disease spread citeepstein2008coupled as our starting point due to its flexibility and ability to produce multiple epidemic waves.
+We take a streamlined version of Epistein et al.’s model of coupled fear and disease spread [21] as our starting point due to its flexibility and ability to produce multiple epidemic waves.
 We then incrementally add the asymptomatic, presymptomatic and exposed disease states present in EpiCast to this model.
 For brevity, we refer to this series of models using the Cartesian product of the two types of states.
-For example, we would refer to a standard model paired with a fear model with Neutral and Afraid states as an coupled model.
+For example, we would refer to a standard model paired with a fear model with Neutral and Fearful states as an coupled model.
 Over the course of this process, we designed the following series of models:
 1.
-, the closest to Epstein et al.’s original model. The flow rate from neutral and afraid compartments is slower by a factor of from the Recovered state. corresponds to the Epstein model.
+, the closest to Epstein et al.’s original model. The flow rate from neutral to fearful compartments is slower by a factor of from the Recovered state. corresponds to the Epstein model.
 2.
-, distinguishing between symptomatic and asymptomatic Infectious and Recovered states. Of the disease states, only has a reduced flow rate from neutral to afraid compartments.
+, distinguishing between symptomatic and asymptomatic Infectious and Recovered states. Of the disease states, only has a reduced flow rate from neutral to fearful compartments.
 3.
@@
 This is the closest to the model implemented in EpiCast.
-For brevity, we present only the system in full. The flow diagram is shown in Fig S1 and the full system of equations is given in Section S1.
+For brevity, we present only the system in full. The flow diagram is shown in Supplementary Fig. S1 and the full system of equations is given in Supplementary Section S1.
 3.2 Coupled contagion agent-based model
@@
 3.2.1 Modeling fear spread
-We implement two methods of fear spread in EpiCast: one which operates based on an agent’s local neighborhood of contacts in a manner similar to disease spread, and a second that approximates the effect of traditional broadcast media.
+[…]
 In order to implement local fear spread, we took advantage of EpiCast’s preexisting methodology for computing disease spread, refactoring the code to allow for fear spread using different per-contact transmission probabilities.
-Fear is spread by symptomatic and afraid agents and countered by agents who know they have recovered from an infection and have a neutral fear state (i.e.
+Fear is spread by symptomatic and fearful agents and countered by agents who know they have recovered from an infection and have a neutral fear state (i.e.
 agents in disease state and fear state ).
@@
 They may either: (1) spread fear, (2) counter fear spread, or (3) remain neutral.
-In order to site these broadcasters, we begin by examining the occupations of agents in each community, which are identified by north American Industry Classification System (NAICS) code in EpiCast.
-If we identify at least one agent with a relevant NAICS code in a given community we locate a broadcaster there and assign all agents with the relevant occupations to work there.
-Each broadcaster then is able to potentially influence anyone within the census tract in which they are located, and starts by taking a neutral position.
+In order to site these broadcasters, we begin by examining the occupations of agents in each community, which are identified by 2017 North American Industry Classification System (NAICS) codes in EpiCast.
+If we identify at least one agent who works in publishing (NAICS code 511) or broadcasting (NAICS code 515) in a given community we locate a media outlet there and assign all agents within the relevant industry to work there.
+Each broadcaster is able to potentially influence anyone within the census tract in which they are located, and starts by taking a neutral position.
 Once initialized, a broadcaster’s position is determined by a combination of the fear states of its employees and the level of disease spread in the broadcast area.
-When the proportion of employees at a broadcaster who are afraid of the disease crosses a certain threshold, , the broadcaster begins spreading fear.
+When the proportion of employees at a broadcaster who are fearful of the disease crosses a certain threshold, , the broadcaster begins spreading fear.
 We then record the current infection rate, the number of new infections in the broadcast area as a proportion of its population, as .
@@
 3.
-Otherwise, the broadcaster spread fear of the disease.
+Otherwise, the broadcaster spreads fear of the disease.
 where .
@@
 3.2.2 Modeling behavior
-We modeled two different ways in which agents may change their behavior when they are afraid: (1) taking protective measures (such as masking and hand-washing) while maintaining their normal routine, and (2) withdrawing from their normal routine and remaining at home.
-We model this first measure by reducing the susceptibility of agents in state , similar to how this is handled in [20] and in the ODE model.
+We modeled two different ways in which agents may change their behavior when they are fearful: (1) taking protective measures (such as masking and hand-washing) while maintaining their normal routine, and (2) withdrawing from their normal routine and remaining at home.
+We model this first measure by reducing the susceptibility of agents in state , similar to how this is handled by Epstein et al.
+[22] and in the ODE model.
 In the second case, we modify an agent’s routine so that they remain in their home community, neither commuting to their work community nor engaging in long distance travel, and only interacting with other members of their household.
@@
 Note that this is different from a complete quarantine, as they can still transmit the disease or their fear to their household.
-If agents are afraid, they can either withdraw purely due to fear with a fixed probability for a given duration , or due to having a symptomatic infection (i.e.
+If agents are fearful, they can either withdraw purely due to fear with a fixed probability for a given duration , or due to having a symptomatic infection (i.e.
 being in state ).
@@
 This was intended to directly address RQ 1 and provide insights into which parameters might be of interest for RQ 3.
-Parameter values are shown in Table S1.
-Second, we sought to compare several scenarios with different types of withdrawals – for hospitalized agents (hospitalized withdrawals), those who are symptomatic and afraid (sick withdrawals), and those who are merely afraid (fear-only withdrawals) – and both purely local (based on physical contact with agents in the same community) and non-local (based on broadcasters operating within a given census tract) fear spread modes in order to address RQ 2.
+Parameter values are shown in Supplementary Table S1.
+Second, we sought to compare several scenarios with different types of withdrawals – for hospitalized agents (hospitalized withdrawals), those who are symptomatic and fearful (sick withdrawals), and those who are merely fearful (fear-only withdrawals) – and both purely local (based on physical contact with agents in the same community) and non-local (based on broadcasters operating within a given census tract) fear spread modes in order to address RQ 2.
 For these scenarios, we use a synthetic population of about 322 million agents representing the 48 contiguous U.S.
@@
 For the first five days of the simulation, for an average of 856 thousand seed cases per day (about 0.265% of the population).
-Parameter values are shown in Table S2.
+Parameter values are shown in Supplementary Table S2.
 Finally, we investigated RQ 3 by conducting a series of sensitivity analyses comparing EpiCast outputs for a range of values of and , which determine the strength of an agent’s behavioral response to fear of the disease.
-We drew these values from the range , trying five values for each parameter.
-We ran each cell twice – once with only local fear spread and once with both local and broadcaster-based fear spread.
+We drew these values from the range , trying five values for each parameter for a total of 25 combinations.
+We ran each combination twice – once with only local fear spread and once with both local and broadcaster-based fear spread.
 These experiments were conducted on the contiguous U.S.
 dataset described above.
-Choices for parameters other than and correspond to Scenarios (c) and (d) from Table S2 (Or, equivalently, Scenarios (e) and (f)) without and with broadcasters, respectively.
+Choices for parameters other than and correspond to Scenarios (c) and (d) from Supplementary Table S2 (or, equivalently, Scenarios (e) and (f)) without and with broadcasters, respectively.
 In order to compute the number of epidemic waves programmatically, we first compute a centered 7-day rolling maximum of new case counts – so as to avoid identifying peaks caused solely by weekly variation – then call the find_peaks function from the Scipy signal package (v1.14.0) in Python with a prominence of .
@@
 For the ODE comparison experiments, Fig.
-1 shows the solutions corresponding the parameter choices that either failed to produce a second epidemic wave within 360 days (top) or produced such a wave (bottom).
-For the model, we observe that changing the likelihood that those who recovered from a symptomatic infection become newly afraid of the disease, , has a dramatic impact on whether or not a second wave occurs.
-With the proportion of the population which are susceptible or afraid largely plateau around 0.8 and 0.6, respectively, after 360 days, as shown in Fig.
+1 shows the solutions corresponding to the parameter choices that either failed to produce a second epidemic wave within 360 days (top) or produced such a wave (bottom).
+For the model, we observe that changing the likelihood that those who recovered from a symptomatic infection become newly fearful of the disease, , has a dramatic impact on whether or not a second wave occurs.
+With the proportion of the population which are susceptible or fearful largely plateau around 80% and 60%, respectively, after 360 days, as shown in Fig.
 1(a).
-In contrast, these proportions plateau at around 0.35 and 0, respectively, when we use , as shown in Fig.
+In contrast, these proportions plateau at around 35% and 0%, respectively, when we use , as shown in Fig.
 1(d).
@@
 When we introduce asymptomatic disease states, as in the model, the situation changes further.
-Keeping all parameters used in the run the same with this model results in 80% of the population still susceptible and 40% afraid after 360 days, and no second epidemic wave, as shown in Fig.
+Keeping all parameters used in the run the same with this model results in 80% of the population still susceptible and 40% fearful after 360 days, and no second epidemic wave, as shown in Fig.
 1(b).
-Changing the susceptibility of afraid agents, , from 25% to 35%, however, is sufficient to produce a second wave in this model, albeit a smaller one which peaks much later at around 275 days, as shown in Fig.
+Changing the susceptibility of fearful agents, , from 25% to 35%, however, is sufficient to produce a second wave in this model, albeit a smaller one which peaks much later at around 275 days, as shown in Fig.
 1(e).
-Under this scenario, about half of the population is still susceptible after 360 days although this proportion is still decreasing.
-The proportion afraid, while approaching zero by 360 days, does so about 160 days later than in the two wave scenario without asymptomatic infections (Fig.
+Under this scenario, about half of the population is still susceptible after 360 days, although this proportion is still decreasing.
+The fearful proportion, while approaching zero by 360 days, does so about 160 days later than in the two wave scenario without asymptomatic infections (Fig.
 1(d)).
@@
 Fig.
-1(c) shows that about 70% and 20% of the population is susceptible and afraid, respectively, after 360 days with in this model.
-Additionally, while we do not see a second wave peak within 360 days here, we do see the start of such a wave and the lack of a plateau of fear levels that corresponds to a single-wave outbreak in the other modes.
+1(c) shows that about 70% and 20% of the population is susceptible and fearful, respectively, after 360 days with in this model.
+Additionally, while we do not see a second wave peak within 360 days here, we do see the start of such a wave and the lack of a plateau of fear levels that corresponds to a single-wave outbreak in the other models.
 When using with the model, Fig.
-1(f) shows that fear levels reach zero and a second epidemic peak as with the model.
+1(f) shows that fear levels reach zero and a second epidemic peak occurs, as with the model.
 However, both these events happen much sooner in the former model, although still about 100 days later than in the model.
@@
 As for RQ 3, we find that changing the values of and both showed the potential to change both whether or not we observe a second wave and the total attack rate.
+Figure 1: ODE model comparison: We show outputs with attention to parameter choices which impact the emergence of multiple epidemic waves.
+On the top and bottom we show solutions corresponding to parameter choices which fail to produce, or produce, respectively, a second wave within 360 days.
+We show these results for the (1(a) and 1(d)), (1(b) and 1(e)), and (1(c) and 1(f)) models.
+For ease of comparison, we combine all infected (including ), recovered, and fearful compartments in each model into a single line in each of the plots above.
 4.3 Comparing selected scenarios in EpiCast
@@
 The scenarios with reduced susceptibility due to fear both buck this trend, with fear levels only falling slightly by day 200 without broadcasters (hosp+sick+reduced_sus) and dropping slowly by around 15 percentage points in total with broadcasters (hosp+sick+reduced_sus+bc).
-We note that, in the absence of broadcasters, the reduced susceptibility case (orange dashed line in Fig.
-2(b)) displays the highest fear levels for most of the simulation and the pure-fear withdrawal case the lowest (red dashed line in Fig.
+We note that, in the absence of broadcasters, the reduced susceptibility case (hosp+sick+reduced_sus in Fig.
+2(b)) displays the highest fear levels for most of the simulation and the pure-fear withdrawal case the lowest (hosp+sick+fear in Fig.
 2(b)).
 In the presence of broadcasters, fear trends more closely resemble other runs.
-Fear levels fall significantly after their peak in the susceptibility reduction case (blue line in Fig.
+Fear levels fall significantly after their peak in the susceptibility reduction case (hosp+sick+reduced_sus+bc in Fig.
 2(b)), when they remaining roughly constant in the absence of broadcasters.
-Conversely, fear levels rose much higher before falling in the pure-fear withdrawal case (green dotted line in Fig.
+Conversely, fear levels rose much higher before falling in the pure-fear withdrawal case (hosp+sick+fear+bc in Fig.
 2(b)), peaking at above 60% of the population rather than less than 40%.
 In both scenarios with broadcasters, more broadcasters spread fear than countered it throughout the majority of the run, though in the pure-fear withdrawal case this changed shortly before the end of the run.
-Far more broadcasters were spreading fear in the susceptibility reduction scenario than in the pure-fear withdrawal scenario for the entire duration (compare blue and green lines in Fig.
+Far more broadcasters were spreading fear in the susceptibility reduction scenario than in the pure-fear withdrawal scenario for the entire duration (compare hosp+sick+reduced_sus+bc and hosp+sick+fear+bc in Fig.
 2(c)), while the reverse is true for fear countering for most of the duration of the run (Fig.
@@
 Broadcasters were not used in the other scenarios.
+Figure 2: EpiCast Fear Spread Scenarios: New cases (2(a)), total fear levels (2(b)), total broadcasters spreading fear (2(c)), and total broadcasters countering fear spread (2(d)) for six different scenarios run in EpiCast.
+These scenarios can include agents withdrawing from their normal schedules due to being hospitalized (hosp), being fearful of the disease while having symptoms of it (sick), and fear of the disease without symptoms (fear), as well as non-local fear spread through broadcast media (bc), and fearful agents having a lower susceptibility to the disease due to taking protective actions such as masking (reduced_sus).
 Fig.
@@
 UT and CO reach similar heights to each other for each peak, while TX has a lower second peak, only slightly above WA.
+Figure 3: Geographic Distribution of Outbreaks in Two Epicast Fear Spread Scenarios: Geographic distribution of new case counts for selected timestamps (left) and overall new case trends for selected states (right) for scenarios with pure-fear withdrawals without (upper) and with (lower) broadcaster-based fear spread.
 4.4 EpiCast sensitivity analysis
 In order to get a better sense of how varying the behavioral response to fear within EpiCast influences the overall trajectory of disease outbreaks, as per RQ 3, we ran a number of simulation varying two influential parameters.
-The first was the likelihood that an agent who is afraid of the disease – but not currently showing symptoms – withdraws from their regular schedule of activities purely due to fear, .
-The second was the scaling factor applied to an afraid agent’s susceptibility to infection, , which represents the impact of afraid agents complying with NPIs that do not directly change their daily schedule, such as masking, maintaining social distance while in public, and hand-washing.
+The first was the likelihood that an agent who is fearful of the disease – but not currently showing symptoms – withdraws from their regular schedule of activities purely due to fear, .
+The second was the scaling factor applied to an fearful agent’s susceptibility to infection, , which represents the impact of fearful agents adopting protective behaviors that do not directly change their daily schedule, such as masking, maintaining social distance while in public, and hand-washing.
 The results of conducting this parameter sweep with only local fear spread are shown in Fig.
@@
 Fig.
-4(a) displays the overall trajectories of each outbreak, with the proportion of new cases shown in orange, that of afraid agents show in blue, and epidemic peaks indicated by the dashed vertical lines.
+4(a) displays the overall trajectories of each outbreak, with the proportion of new cases shown in orange, that of fearful agents show in blue, and epidemic peaks indicated by the dashed vertical lines.
 Note that all parameter combinations produce a single epidemic wave.
@@
 The lowest attack rates, in contrast, are observed when and are both small.
-Higher values of correspond to fewer afraid agents and result reduced variation in fear trajectories for different values of .
+Higher values of correspond to fewer fearful agents and result reduced variation in fear trajectories for different values of .
 For , we observe that fear levels rapidly plateau for and rise to a peak before falling to varying degrees for .
-This drop in fear levels is more pronounced for greater values of , when the overall attack rate is greater.
+This drop in fear levels is more pronounced for larger values of , when the overall attack rate is higher.
+Figure 4: Epicast local fear spread sensitivity analysis: New cases by day (with peaks indicated by dashed vertical line) for (4(a)) a coarse-grained parameter search and (4(b)) a finer-grained search highlighting cases with multiple waves, and (4(c)) total attack rate for various rates of purely fear-based withdrawals () and levels of susceptibility scaling due to fear () with only local fear spread.
 The results of conducting this parameter sweep with both local and broadcaster-based fear spread are shown in Fig.
@@
 The overall trajectories of each outbreak are shown in Fig.
-5(a) with the dashed vertical lines indicating where peaks occur, and the orange and blue lines indicating proportion of new cases and afraid agents, respectively.
+5(a) with the dashed vertical lines indicating where peaks occur, and the orange and blue lines indicating proportion of new cases and fearful agents, respectively.
 Fig.
@@
 The height of the initial peaks decreases as increases and fear levels fall farther when more cases occur.
+Figure 5: Epicast broadcaster fear spread sensitivity analysis: (5(a)) new cases by day (with peaks indicated by dashed vertical line), (5(b)) number of epidemic waves, and (5(c)) total attack rate for various rates of purely fear-based withdrawals () and levels of susceptibility scaling due to fear () with both local and broadcaster-based fear spread.
 5 Discussion
@@
 2(b), offers some insights into why we might observe multiple peaks primarily in cases with non-local fear spread via broadcasters.
-In cases where fear reduces agent susceptibility to infection, in the absence of fear spread through broadcasters (hosp+sick+reduced_sus) fear levels never fall far below their peak around 50 days.
-This is likely due to infection levels never rising to the point where there is a sufficient number of agents in the state to cause afraid agents to lose their fear.
+In cases where fear reduces agent susceptibility to infection, in the absence of fear spread through broadcasters (hosp+sick+reduced_sus) fear levels never fall far below their peak, at around 50 days.
+This is likely due to infection levels never rising to the point where there is a sufficient number of agents in the state to cause fearful agents to lose their fear.
 This is in contrast to what occurs with fear spread through broadcasters (hosp+sick+reduced_sus+bc), where fear levels quickly fall after reaching their peak.
@@
 Roughly the opposite appears to occur in the cases where we allow purely fear-based withdrawals; without the influence of broadcasters (hosp+sick+fear), fear levels remain lower than in any other scenario for the duration of the run.
-This is likely because withdrawals reduce both fear and disease spread in in this mode there is a high chance that any given afraid person will withdraw.
+This is likely because withdrawals reduce both fear and disease spread in this mode and there is a high chance that any given fearful person will withdraw.
 Conversely, with the influence of broadcasters (hosp+sick+fear+bc) fear levels rise much higher before falling, though they remain lower than in all other scenarios except that with only hospitalization-based withdrawals.
@@
 2(d)) over time.
-The period from about 25-50 where fear levels peak and then begin to fall in the reduced susceptibility scenario with broadcasters is also the period when we see the number of broadcasters countering fear spread dramatically increase before reaching a much lower sustained growth rate which again corresponds to a slow linear decrease in the number of afraid agents.
-The number of fear spreading broadcasters follows a similar trend to fear levels, here, albeit with a smaller falloff after the peak.
-Similarly, we see overall fear levels and the number of fear-spreading broadcasters in the pure-fear withdrawal case, with the dramatic reduction in fearful agents between days 50 and 100 matched by decline in fear-spreading broadcasters of similar magnitude.
+The period from days 25-50 where fear levels peak and then begin to fall in the reduced susceptibility scenario with broadcasters is also the period when we see the number of broadcasters countering fear spread dramatically increase, before reaching a much slower sustained growth rate.
+This slower growth period corresponds to a slow linear decrease in the number of fearful agents.
+The number of fear spreading broadcasters follows a similar trend to fear levels in this case, albeit with a smaller falloff after the peak.
+Similarly, we see that overall fear levels and the number of fear-spreading broadcasters show similar patterns in the pure-fear withdrawal case, with the dramatic reduction in fearful agents between days 50 and 100 matched by decline in fear-spreading broadcasters of similar magnitude.
 In addition, the very narrow range of parameter values for which we do observe multiple waves with purely local fear spread appears to fall in the midst of a phase transition between a regime of small, short outbreaks and large, long outbreaks as represented by the cell , and the cell in Fig.
-4(a).
+4(a), respectively.
 The former regime corresponds to lower values of which represent a strong NPI response to fear, implying that fear-induced behavior reduces disease spread sufficiently to forestall a larger outbreak here.
@@
 In other words, in order to produce multiple epidemic waves in this model, fear must first spread significantly faster than the disease and then recede faster than it as well.
-Within this framing, it is clear why broadcasters facilitate the formation of multiple waves: information (and the behavioral changes it engenders) both spreads and is forgotten more rapidly when it is less tightly constrained by physical distance.
+Within this framing, it is clear why broadcasters facilitate the formation of multiple waves: information and the behavioral changes it engenders both spreads and recedes more rapidly when it is less tightly constrained by physical distance.
 The results we present in this paper are perhaps best understood as an exploration of the high-level implications of our model.
 Though the foundational elements of EpiCast are mature and have been calibrated extensively, there remain significant obstacles to doing so for the fear-based components of the model.
-While some survey data is available describing various behaviors throughout part of the pandemic [27], it is more difficult to tie specific behaviors to specific influences or rationales.
+While some survey data is available describing various behaviors throughout part of the pandemic [29], it is much more difficult to tie behavior to specific influences or rationales.
 We thus confine ourselves to presenting the model itself and exploring the impact of varying parameter values, and leave calibration efforts to future work.
@@
 In addition, there are many simulation parameters beyond the two we chose to study.
-Extending the parameter sweep method to the other model parameters would be prohibitively expensive, but in the absence of such an experiment it is unclear how varying those parameters might affect the dynamics in Epicast.
+Extending the parameter sweep method to the other model parameters would be prohibitively computationally expensive, but in the absence of such an experiment it is unclear how varying those parameters might affect the dynamics in EpiCast.
 6 Conclusion
-In this work we presented a series of models which capture the coupled spread of fear of a disease and the disease itself.
+In this work we present a series of models which capture the coupled spread of fear of a disease and the disease itself.
 We do so with several ODE models in order to demonstrate the impact of adding additional disease states to the model, representing asymptomatic, exposed, and pre-symptomatic disease states, and an ABM covering all disease states with both local and non-local mechanisms for fear spread implemented.
 We find that the total attack rate is reduced by the addition of asymptomatic states and increased by the addition of a pre-symptomatic state in our ODE models, and the reverse is true for the rate at which an outbreak progresses.
-In our ABM implementation in EpiCast, we primarily observe multiple epidemic waves when using our non-local fear spread mechanism, which represents traditional broadcast media.
-We find that a combination of (1) high initial fear levels, (2) a strong behavioral response to fear, and (3) low fear levels later in the simulation produces multiple waves in our model.
+In our ABM implementation in EpiCast, we primarily observe multiple epidemic waves when using our non-local fear spread mechanism, which represents broadcast media (regional information communication, including print, radio, and TV).
+We find that a combination of (1) high initial fear levels, (2) a strong behavioral response to fear, and (3) low fear levels later in the simulation produce multiple waves in our model.
 Thus, one potential explanation for why we only observe multiple waves with purely local fear spread for a very small range of parameter choices is that fear generally spreads (and recedes) much more slowly in the absence of broadcasters in our model.
 Acknowledgments
-This work was supported by the U.S.
-Department of Energy, Office of Science, Office of Advanced Scientific Computing Research and by cooperative agreement CDC-RFA-FT-23-0069 from the CDC’s Center for Forecasting and Outbreak Analytics.
-Its contents are solely the responsibility of the authors and do not necessarily represent the official views of the Centers for Disease Control and Prevention.
-This work was performed at Los Alamos National Laboratory (LANL), an equal opportunity employer, which is operated by Triad National Security, LLC, for the National Nuclear Security Administration (NNSA) of the U.S.
-Department of Energy (DOE) under contract #19FED1916814CKC.
-The funders had no role in study design, data collection and analysis, decision to publish, or preparation of the manuscript.
-The findings and conclusions in this report are those of the authors and do not necessarily represent the official position of LANL.
-This research used resources provided by the Darwin testbed at LANL which is funded by the Computational Systems and Software Environments subprogram of LANL’s Advanced Simulation and Computing program (NNSA/DOE).
-This material is based upon work supported by the U.S.
-Department of Energy, Office of Science, Office of Advanced Scientific Computing Research, Department of Energy Computational Science Graduate Fellowship under Award Number DE-SC0021
+This work was supported by the U.S. Department of Energy,
+Office of Science, Office of Advanced Scientific Computing
+Research and by cooperative agreement CDC-RFA-FT-23-0069 from
+the CDC’s Center for Forecasting and Outbreak Analytics.
+Its contents are solely the responsibility of the authors and
+do not necessarily represent the official views of the Centers
+for Disease Control and Prevention.
+This work was performed at
+Los Alamos National Laboratory (LANL), an equal opportunity
+employer, which is operated by Triad National Security, LLC,
+for the National Nuclear Security Administration
+(NNSA) of the U.S. Department of Energy (DOE) under contract
+#19FED1916814CKC. The funders had no role in study design,
+data collection and analysis, decision to publish, or preparation
+of the manuscript. The findings and conclusions in this report
+are those of the authors and do not necessarily represent
+the official position of LANL.
+This research used resources provided by the Darwin testbed at
+LANL which is funded by the Computational Systems and Software
+Environments subprogram of LANL’s Advanced Simulation and
+Computing program (NNSA/DOE).
+This material is based upon work supported by the U.S. Department
+of Energy, Office of Science, Office of Advanced Scientific
+Computing Research, Department of Energy Computational Science
+Graduate Fellowship under Award Number DE-SC0021.
 Author contributions statement
-J.K., P.C.A, S.Y.V., and T.C.G.
+J.K., P.C.A, D.J.B., C.B., S.Y.D.V., and T.C.G.
 developed the conceptual framework and J.K.
@@
 contributing some code to the later.
+D.J.B.
+and C.B.
+provided suggestions regarding the development and evaluation of the ODE models.
 P.C.A.
@@
 provided insights into the original design and implementation of EpiCast which informed the implementation of the behavioral model.
-J.K., A.B., S.Y.V., and T.C.G.
+J.T.
+developed and provided access to UrbanPop, the synthetic population on which this study was conducted, and guided that use.
+J.K., A.B., S.Y.D.V., and T.C.G.
 conceived the experiments, and J.K.
@@
 prepared the initial draft of the manuscript, along with all figures.
-All authors reviewed the manuscript.
+A.B., S.Y.D.V., and T.C.G.
+all provided supervision.
+All authors reviewed and edited the manuscript.
 Data availability statement
-We intend to release all output data from the simulations described in this paper available via Zenodo link upon publication, along with all scripts necessary to generate the plots contained within this manuscript from those results.
-These scripts are unfortunately still in the process of being reviewed for release by the project sponsors and are thus not available at this time.
+We have publicly released all output data from simulations described in this paper through Zenodo here: https://zenodo.org/uploads/15706577.
+The source code for the scripts used to analyze these results and produce the plots shown in this manuscript are publicly available here: https://github.com/lanl/epicast-coupled.
+While we hope to eventually release the full code of both EpiCast and the fear spread extension presented in this manuscript, this process is especially long for patentable code given the policies of Los Alamos National Laboratory (where EpiCast is developed) and the multiple U.S.
+government agencies which sponsored this work.
+Limited access to this code may be granted upon request.
 Competing interests
 The authors declare no competing interests.
-Supplementary Material
-S1 Coupled contagions ODE models
-We provide the full formulation of the system of ODEs which define the model below.
-This system is summarized in the flow diagram in Fig.
-S1.
-Let be the size of the compartment with disease state and fear state .
-The system is as follows:
 S2 Model parameters
-Parameter values used in ODE model comparison experiments (Section 4.2) are shown in Table S1 while those used in the EpiCast scenario comparision experiments (Section 4.3) are shown in Table S2.
-Variable | Description | Experiment | ||||||||
-(a) | (b) | (c) | (d) | (e) | (f) | |||||
-disease transmission rate | 0.2 | 0.2 | 0.2 | 0.2 | 0.2 | 0.2 | ||||
-disease recovery rate | 0.5 | 0.5 | 0.5 | 0.5 | 0.5 | 0.5 | ||||
-0.6 | 0.6 | 0.6 | 0.6 | ||||||
-0.4 | 0.4 | 0.4 | 0.4 | ||||||
-0.75 | 0.75 | 0.75 | 0.75 | ||||||
-0.5 | 0.5 | ||||||||
-0.22 | 0.22 | 0.22 | 0.22 | 0.22 | 0.22 | ||||
-baseline fear loss rate | 0.05 | 0.05 | 0.05 | 0.05 | 0.05 | 0.05 | ||||
-fear loss contact rate | 0.44 | 0.44 | 0.44 | 0.44 | 0.44 | 0.44 | ||||
-1 | 0 | 0 | 0 | 0 | 0 | ||||
-1 | 1 | 1 | 1 | 1 | 1 | ||||
-0.25 | 0.25 | 0.25 | 0.35 | 0.25 | 0.35 |
-Variable | Description | Scenario | ||||||||
-(a) | (b) | (c) | (d) | (e) | (f) | |||||
-disease transmission rate | 0.2 | 0.2 | 0.2 | 0.2 | 0.2 | 0.2 | ||||
-disease recovery rate | 0.5 | 0.5 | 0.5 | 0.5 | 0.5 | 0.5 | ||||
-0.6 | 0.6 | 0.6 | 0.6 | 0.6 | 0.6 | ||||
-0.4 | 0.4 | 0.4 | 0.4 | 0.4 | 0.4 | ||||
-0.75 | 0.75 | 0.75 | 0.75 | 0.75 | 0.75 | ||||
-0.5 | 0.5 | 0.5 | 0.5 | 0.5 | 0.5 | ||||
-0.22 | 0.22 | 0.22 | 0.22 | 0.22 | 0.22 | ||||
-baseline fear loss rate | 0.05 | 0.05 | 0.05 | 0.05 | 0.05 | 0.05 | ||||
-fear loss contact rate | 0.44 | 0.44 | 0.44 | 0.44 | 0.44 | 0.44 | ||||
-0 | 0 | 0 | 0 | 0 | 0 | ||||
-1 | 1 | 1 | 1 | 1 | 1 | ||||
-1 | 1 | 1 | 1 | 0.35 | 0.35 | ||||
-0 | 1 | 1 | 1 | 1 | 1 | ||||
-0 | 0 | 0.65 | 0.65 | 0 | 0 | ||||
-0 | 0 | 0 | 0.25 | 0 | 0.25 | ||||
-0.5 | 0.5 | ||||||||
-0.25 | 0.25 | ||||||||
-0.125 | 0.125 |
+Parameter values used in ODE model comparison experiments (Section 4.2) are shown in Suplementary Table S1 while those used in the EpiCast scenario comparision experiments (Section 4.3) are shown in Suplementary Table S2.
+Variable
+Description
+Experiment
+disease transmission rate
+0.2
+0.2
+0.2
+0.2
+0.2
+0.2
+disease recovery rate
+0.5
+0.5
+0.5
+0.5
+0.5
+0.5
+probability of
+symptomatic infection
+0.6
+0.6
+0.6
+0.6
+probability of
+asymptomatic infection
+0.4
+0.4
+0.4
+0.4
+relative infectivity of
+asymptomatic infection
+0.75
+0.75
+0.75
+0.75
+inverse of
+incubation period
+0.5
+0.5
+fear transmission
+rate
+0.22
+0.22
+0.22
+0.22
+0.22
+0.22
+baseline fear loss rate
+0.05
+0.05
+0.05
+0.05
+0.05
+0.05
+fear loss contact rate
+0.44
+0.44
+0.44
+0.44
+0.44
+0.44
+relative fear
+susceptibility after
+symptomatic recovery
+1
+0
+0
+0
+0
+0
+relative infectivity
+when fearful
+1
+1
+1
+1
+1
+1
+relative susceptibility
+when fearful
+0.25
+0.25
+0.25
+0.35
+0.25
+0.35
+Table S1: Parameter values used in ODE model experiments: Results of experiments (a-e) are shown in the corresponding subfigures in Fig.
+1, and the EpiCast scenarios to setups with (a) only hospitalization-based withdrawals, (b) the addition of symptomatic fearful withdrawals, and the addition of either pure-fear withdrawals from fear – without (c) or with (d) broadcaster-based fear spread – or reduced susceptibility when fearful – without (e) or with (f) broadcaster-based fear spread.
+Variable
+Description
+Scenario
+disease transmission rate
+0.2
+0.2
+0.2
+0.2
+0.2
+0.2
+disease recovery rate
+0.5
+0.5
+0.5
+0.5
+0.5
+0.5
+probability of
+symptomatic infection
+0.6
+0.6
+0.6
+0.6
+0.6
+0.6
+probability of
+asymptomatic infection
+0.4
+0.4
+0.4
+0.4
+0.4
+0.4
+relative infectivity of
+asymptomatic infection
+0.75
+0.75
+0.75
+0.75
+0.75
+0.75
+inverse of
+incubation period
+0.5
+0.5
+0.5
+0.5
+0.5
+0.5
+fear transmission
+rate
+0.22
+0.22
+0.22
+0.22
+0.22
+0.22
+baseline fear loss rate
+0.05
+0.05
+0.05
+0.05
+0.05
+0.05
+fear loss contact rate
+0.44
+0.44
+0.44
+0.44
+0.44
+0.44
+relative fear
+susceptibility after
+symptomatic recovery
+0
+0
+0
+0
+0
+0
+relative infectivity
+when fearful
+1
+1
+1
+1
+1
+1
+relative susceptibility
+when fearful
+1
+1
+1
+1
+0.35
+0.35
+probability of withdrawal
+when fearful with symptoms
+0
+1
+1
+1
+1
+1
+probability of pure-fear
+withdrawal when fearful
+0
+0
+0.65
+0.65
+0
+0
+probability of watching
+a given broadcaster
+0
+0
+0
+0.25
+0
+0.25
+threshold of fearful
+workers for broadcasters
+to take position
+0.5
+0.5
+relative rate of new cases
+for broadcasters to
+resume neutral position
+0.25
+0.25
+relative rate of new cases
+for broadcasters
+to counter fear spread
+0.125
+0.125
+Table S2: Parameter values used in EpiCast scenario experiments: Parameters represent scenarios with (a) only hospitalization-based withdrawals, (b) the addition of symptomatic fearful withdrawals, and the addition of either pure-fear withdrawals – without (c) or with (d) broadcaster-based fear spread – or reduced susceptibility when fearful – without (e) or with (f) broadcaster-based fear spread.